What is the pH of a solution that results from diluting 0.700 mol acetic acid (CH3CO2H) and 0.129 mol sodium acetate (NaCH3CO2) with water to a volume of 1.41 L? (Ka of CH3CO2H = 1.8E-5)
Answer #1

Ka = 1.8*10^-5

pKa = - log (Ka)

= - log(1.8*10^-5)

= 4.745

use:

pH = pKa + log {[conjugate base]/[acid]}

Since volume is same for numerator and denominator, above expression can also be written as:

pH = pKa + log {mol(conjugate base)/mol(acid)}

= 4.745+ log {0.129/0.7}

= 4.01

Answer: 4.01